# Instant Runoff Voting and the Reinforcement Paradox

### v2

@@ 6. Empirical Results
In total, we accessed the preference profiles for 352 single-winner IRV elections with at least three candidates. [^4] Any reference to the number of candidates in an election does not include write-ins. In our dataset, write-in candidates receive an insignificant amount of votes and ignoring them does not affect the analysis. Many of these elections contain more than three candidates; because we are interested in the three-candidate case, for each election we “reduce” the number of candidates by running the IRV algorithm until only three candidates remain and then analyze the resulting profile. Note that the relationship between the original election and the resulting three-candidate election is not clear with respect to the susceptibility to reinforcement paradoxes. It is a priori possible that the original election demonstrates a paradox while the reduced three-candidate election does not, and vice versa. We use this methodology because we are interested in three-candidate elections and analyzing the elections with more than three candidates in this way the IRV algorithm in this allows us to expand the size of our three-candidate election dataset. Moreover, in some cases if the resulting three-candidate election demonstrates a paradox then we can gain insight into how the original election demonstrates a paradox; this is how a reinforcement paradox in a five-candidate Minneapolis city council election was found, for example [25].
@@ 6.2. Bootstrapping Results
In Section 5 we ran simulations based on sampling using theoretical models such as the IAC model. Since we have access to a large dataset of real-world elections, we can also sample from these elections and run simulations using nonparametric bootstrap analysis. To perform this analysis, we follow techniques used in the classical papers [38] and [40]. For each of the 352 three-candidate profiles we have from our real-world dataset, we created 10,000 pseudoprofiles by sampling ballots with replacement from the actual three-candidate profile. This sampling allowed us to generate 3,520,000 three-candidate elections which we can check for reinforcement paradoxes. To keep the computation time reasonable and to match the analysis in Section 5, if the actual profile contained more than 1,001 voters then for each generated pseudoprofile we sampled only 1,001 ballots; otherwise each pseudoprofile contained the same number of voters as the original election. Most of the elections in the dataset contain more than 1,001 voters and thus most of the generated elections contain 1,001 voters.
@@
However, bootstrapping analysis can give a sense of how “close” a given election comes to demonstrating a paradox. If none of the 10,000 pseudoprofiles generated from an election demonstrate a paradox then the original election is not “close” to an election which demonstrates a paradox. On the other hand, if an election demonstrates a paradox but relatively few of its pseudoprofiles do, then the election is close to one which does not demonstrate a paradox. For example, the 2023 city council ward 7 election in Minneapolis (Table 16) demonstrated a reinforcement paradox but only 3370 of 10000 generated pseudoprofiles demonstrate a paradox. Thus, we could argue that this election demonstrating a paradox is not the most likely outcome.
@@
A more nuanced summary of the bootstrapping results can be seen in Figure 3. Each histogram illustrates the frequency of the proportion of pseudoprofiles generated by each original election that resulted in a “paradoxical” election. The top image is based on the 86 original three-candidate elections that demonsrated the paradox. Thus, for example, the right-most bar on the top histogram indicates that in 33 of these elections, between 95 and 100% of the associated pseudoprofiles exhibited the paradox. Similarly, the second-to-the-right bar indicates that in 10 of these elections, between 90 and 95% of the associated pseudoprofiles exhibited the paradox. The bottom histogram is constructed the same way for the 266 original three-candidate elections which did not exhibit the paradox with the exception that elections which generated zero “paradoxical” pseudoprofiles are omitted for clarity (since the left bar would be out-sized).
@@
From Figure 3 (top), we can observe that roughly half of the original elections exhibiting a paradox were extremely likely to exhibit that paradox in the sense that more than 90% of their associated pseudoprofiles also did so. Most of the remaining elections in this category also generated pseudoprofiles that more often than not exhibited a reinforcement paradox. The reverse is true in Figure 3 (bottom): the large majority of the original elections that did not exhibit the paradox generated few pseudoprofiles exhibiting a reinforecement paradox.
@@
Noteworthy, however, is the fact that across all 352 original elections, 206 generated at least one pseudoprofile which demonstrated a reinforcement paradox. Thus, arguably, 206 elections are close to an election which exhibits a paradox. If we restrict the analysis to elections that did not contain a majority candidate, 187 of the 208 of the original elections generated at least one pseudoprofile exhibiting a reinforcement paradox. These results suggest that almost any three-candidate real-world election which does not contain a majority candidate is “close” to an election demonstrating a reinforcement paradox. We note such a high percentage is not achieved under bootstrapping for other paradoxes. For example, in [27] the same type of analysis was performed for the spoiler effect using a smaller dataset of American IRV elections (but still using 10,000 pseudoprofiles per election), and only five elections generated at least one pseudoprofile demonstrating a spoiler effect under IRV. Thus, it is not the case that performing this bootstrap analysis guarantees that almost all elections in a dataset will be “close” to an election which exhibits a given paradox.
@@
Recall in the 352 actual elections, any election demonstrating a plurality loser paradox also demonstrated a runner-up paradox. This is not the case for pseudoprofiles: 9,542 generated elections demonstrate a plurality loser but not a runner-up paradox. Thus, while such an outcome is not likely, some real-world elections are close to this phenomenon occurring.
@@
Figure 3. (Top) Bootstrapping frequency results for elections which demonstrated a reinforcement paradox. (Bottom) Bootstrapping frequency results for elections which did not demonstrate a reinforcement paradox but generated at least one pseudoprofile that did.
@@ 7. Discussion
In this section we discuss our results, starting with a comparison between the theoretical and empirical results, then placing our results into context by comparing them to other classical paradoxes, and concluding by comparing our results to previous results on the reinforcement paradox.
@@ 7.1. Theoretical Versus Empirical Results
As is often the case in social choice, none of the theoretical models give probabilities which closely match what we observe in real-world elections across all four paradox types. Interestingly, if we focus on the 352 three-candidate reduced elections (lines two and four of Table 15), the IAC complete-ballot model (Table LABEL:IAC_results) provides a decent approximation of the empirical results. Furthermore, even though the real-world elections contain many partial ballots, generally the complete-ballot IAC model more accurately reflects the empirical probabilities across the 352 reduced elections than the partial-ballot IAC model. For other voting paradoxes such as monotonicity paradoxes or the spoiler effect, the IAC model typically gives probabilities much higher than what we observe in real elections. If we focus only on the 151 three-candidate actual elections, Dir$(0.6)$ gives the best approximation of the empirical unconditioned results (with the exception of “PL and RU”), which makes sense because in these elections, the winner is stronger overall than in the Dir$(1)$ (IAC) model, consistent with the election data. Overall, the 1D-spatial models, both normal and bimodal, underestimate the probabilities of one or more types of paradox occurring.
@@
7.2. Comparison to Other Paradoxes
Compared to other classical paradoxes such as downward and upward monotonicity paradoxes, no-show paradoxes, or the spoiler effect, the probability an election demonstrates a reinforcement paradox is much larger. For example, Graham-Squire and McCune [17] use a very similar database to provide empirical probabilities, conditioned on the absence of a majority candidate, of 2.2%, 1.6%, 0.5%, and 2.2% for upward monotonicity paradoxes, downward monotonicity paradoxes, no-show paradoxes, and the spoiler effect, respectively. The unconditioned probabilities fall to around 1% or less. Theoretical probabilities for these other paradoxes under models like IAC or IC also tend to be much lower than what we observe for reinforcement paradoxes [18, 27]. In particular, we are not aware of any other paradox for which the IC model returns a probability of 1.
As far as we are aware, prior to this work the paradox which gave the highest probabilities is the negative participation paradox, explored in [18, 28]. McCune and Wilson [28] give an unconditioned empirical probability of 3.3%, which rises to 4.9% when we condition on the absence of a majority candidate. They also show that for a very large electorate with complete ballots, the unconditioned probability of observing this paradox is 15.77% under the IC model, and Kamwa et al. [18] give a probability of 7.69% under the IAC model. These probabilities are significantly smaller than the reinforcement paradox probabilities.
@@ 7.3. Comparison to Previous Reinforcement Paradox Results
As mentioned in the Introduction, we are aware of only two previous papers [8, 37] which estimate the probability with which an election demonstrates a reinforcement paradox. Our work differs from theirs in several significant ways. Most importantly, in most cases, the analysis in these earlier papers was based on simulating two separate elections, checking whether the elections had the same winner, and then indicating a reinforcement paradox had occurred if the winner obtained from the combined election was different from the winner in the two generated elections. That is, the authors did not start with one election and look for a partition of the ballots as we do here. As Plassmann and Tideman state in [37], “We determine the frequencies of the reinforcement paradox by sampling ballots for two separate elections, which we then combine to determine whether the paradox occurred for any voting rule. In contrast, we determine the frequencies of all other paradoxes from a single sample of ballots.” Plassmann and Tideman perhaps made this choice from convenience, as they later state, “There is no obvious way in which to divide the ballots from a single election into two distinct sets that would yield the appropriate distribution of rankings in the two new sets and would permit us to examine reinforcement.” We have contributed to their work by providing such a way to divide the ballots in the 3-candidate case in Propositions 1 and 2. Courtin et al. [8] generally take a similar approach, simulating two ballot sets and then combining them. We feel our approach is equally valid since we are interested in the “paradoxical harm” felt by a candidate (and perhaps their supporters) in an election exhibiting the paradox: if there exists a partition of the electorate such that one of the losing candidates can point to each sub-electorate and claim victory, that candidate can reasonably claim some sort of harm when they ultimately lose the overall election.
@@
Finally, both previous papers report extremely small probabilities for observing a reinforcement paradox under IRV. Plassmann and Tideman give a probability of approximately 0.02% when using an electorate of at least 1,000 voters. Courtin et al. report similarly small probabilities. Using the term “Young’s reinforcement axiom” (YRA) synonymously with Young’s notion of consistency (Definition 1), Courtin et al. state, “Our results show that … cases of violations of YRA are rather rare… [[[[A]]]]lthough all these [[[[iterative scoring]]]] rules violate YRA, the frequencies of violation are very rare. Then, since the violation of this axiom is not so frequent, one should not be unduly worried about its theoretical possibility.” By contrast, we report extremely large reinforcement paradox probabilities, which are generally much larger than the probabilities of observing other classical paradoxes. The reason for this discrepancy is our difference in approach, as outlined above. We would also obtain similarly small probabilities if we were to take each real election and search for reinforcement paradoxes by choosing a partition of the ballots at random, for example. We obtain higher probabilities because we control how the ballots are partitioned.
@@ 8. Conclusion
Both theoretical and empirical analysis suggests that the reinforcement paradox is significantly more likely to occur under IRV with 3 candidates than other paradoxes. This is partly because, as noted above, there are no restrictions on how the ballots are partitioned. If we require the two sub-elections to be similarly-sized, for example, as they might be in two precincts, there will most likely be fewer instances of the paradox; the same is true if we restrict the sub-elections so that one sub-election is a small perturbation of the original election. Future work could examine how the probability of a reinforcement paradox changes under restrictions of this kind, or to instances involving a larger number of sub-elections. Finally, this work could be extended to more than three candidates by determining conditions on the parameters in an $n$-candidate preference profile under which the paradox occurs.
The results in this paper also shed light on ways in which IRV differs from positional scoring methods, which are not susceptible to the reinforcement paradox. It would be interesting to repeat this analysis for other voting methods, such as Condorcet consistent methods, that do not fall into this class.

### v3

@@ 5. Empirical Results
In total, we accessed the preference profiles for 352 single-winner IRV elections with at least three candidates. [^3] Any reference to the number of candidates in an election does not include write-ins. In our dataset, write-in candidates receive an insignificant amount of votes and ignoring them does not affect the analysis. Many of these elections contain more than three candidates; because we are interested in the three-candidate case, for each election we “reduce” the number of candidates by running the IRV algorithm until only three candidates remain and then analyze the resulting profile. Note that the relationship between the original election and the resulting three-candidate election is not clear with respect to the susceptibility to reinforcement paradoxes. It is a priori possible that the original election demonstrates a paradox while the reduced three-candidate election does not, and vice versa. We use this methodology because we are interested in three-candidate elections and analyzing the elections with more than three candidates using the IRV algorithm in this way allows us to expand the size of our three-candidate election dataset. Moreover, in some cases if the resulting three-candidate is susceptible to a paradox then we can gain insight into how the original election might be susceptible to a paradox; this is how a reinforcement paradox in a five-candidate Minneapolis city council election was found, for example [35].
@@ 5.2. Bootstrapping Results
In Section 4 we ran simulations based on sampling using theoretical models such as the IAC model. Since we have access to a large dataset of real-world elections, we can also sample from these elections and run simulations using nonparametric bootstrap analysis.
Bootstrapping allows us to generate random samples from each election profile to create a set of “pseudoprofiles.” Analysis of these pseudoprofiles can give a sense of how “close” the given election is to demonstrating a paradox. For instance, if we sample from an election profile that does not demonstrate a reinforcement paradox and determine that none of the sampled pseudoprofiles demonstrates a paradox, then the original election is not close to an election that demonstrates a paradox. On the other hand, if an election demonstrates a paradox but relatively few of its pseudoprofiles do, then the election is close to one which does not demonstrate a paradox. This is the case, for example, in the 2023 city council ward 7 election in Minneapolis (Table 16). The election demonstrates a reinforcement paradox but only 3370 of 10000 generated pseudoprofiles demonstrate a paradox. Thus, we could argue that this election demonstrating a paradox is not the most likely outcome.
To perform this analysis, we follow techniques used in the classical papers [48] and [50]. For each of the 352 three-candidate profiles from our real-world dataset, we created 10,000 pseudoprofiles by sampling ballots with replacement from the actual three-candidate profile. This sampling allowed us to generate 3,520,000 three-candidate elections which we can check for reinforcement paradoxes.
To keep the computation time reasonable and to match the analysis in Section 4, if the actual profile contained more than 1,001 voters then for each generated pseudoprofile we sampled only 1,001 ballots; otherwise each pseudoprofile contained the same number of voters as the original election. Most of the elections in the dataset contain more than 1,001 voters and thus most of the generated elections contain 1,001 voters.
@@
A more nuanced summary of the bootstrapping results can be seen in Figure 3. Each histogram illustrates the frequency of the proportion of pseudoprofiles generated by each original election that resulted in a “paradoxical” election. The top image is based on the 86 original three-candidate elections that were susceptible to the paradox. Thus, for example, the right-most bar on the top histogram indicates that in 33 of these elections, between 95 and 100% of the associated pseudoprofiles were susceptible to the paradox. Similarly, the second-to-the-right bar indicates that in 10 of these elections, between 90 and 95% of the associated pseudoprofiles were susceptible to the paradox. The bottom histogram is constructed the same way for the 266 original three-candidate elections which were not susceptible to the paradox with the exception that elections that generated zero “paradoxical” pseudoprofiles are omitted for clarity (since the left bar would be out-sized).
@@
From Figure 3 (top), we can observe that roughly half of the original elections that were susceptible to a paradox were extremely likely to exhibit that paradox in the sense that more than 90% of their associated pseudoprofiles also did so. Most of the remaining elections in this category also generated pseudoprofiles that more often than not were susceptible to a reinforcement paradox. The reverse is true in Figure 3 (bottom): the large majority of the original elections that were not susceptible to the paradox generated few pseudoprofiles that were susceptible to a reinforcement paradox.
@@
Noteworthy, however, is the fact that across all 352 original elections, 206 generated at least one pseudoprofile which was susceptible to a reinforcement paradox. Thus, arguably, 206 elections are close to an election that is susceptible to a paradox. If we restrict the analysis to elections that did not contain a majority candidate, 187 of the 208 of the original elections generated at least one pseudoprofile that is susceptible to a reinforcement paradox. These results suggest that almost any three-candidate real-world election which does not contain a majority candidate is “close” to an election that is susceptible to a reinforcement paradox. We note such a high percentage is not achieved under bootstrapping for other paradoxes. For example, in [37] the same type of analysis was performed for the spoiler effect using a smaller dataset of American IRV elections (but still using 10,000 pseudoprofiles per election), and only five elections generated at least one pseudoprofile demonstrating a spoiler effect under IRV. Thus, it is not the case that performing bootstrap analysis guarantees that almost all elections in a dataset will be “close” to an election that exhibits a given paradox.
@@
Finally, recall in the 352 actual elections, any election that was susceptible to a plurality loser paradox was also susceptible to a runner-up paradox. This is not the case for pseudoprofiles: 9,542 generated elections are susceptible to a plurality loser but not to a runner-up paradox. Thus, while such an outcome is not likely, some real-world elections are close to this phenomenon occurring.
@@
Figure 3. (Top) Bootstrapping frequency results for elections that are susceptible to a reinforcement paradox. (Bottom) Bootstrapping frequency results for elections that are are not susceptible to a reinforcement paradox but generated at least one pseudoprofile that is.
@@ 6. Discussion
In this section we discuss our results, starting with a comparison between the theoretical and empirical results, then placing our results into context by comparing them to previous results on the reinforcement paradox.
@@ 6.1. Theoretical Versus Empirical Results
As is often the case in social choice, none of the theoretical models give probabilities which closely match what we observe in real-world elections across all four paradox types. Interestingly, if we focus on the 352 three-candidate reduced elections (lines two and four of Table 15), the IAC complete-ballot model (Table 11) provides a decent approximation of the empirical results. Furthermore, even though the real-world elections contain many partial ballots, generally the complete-ballot IAC model more accurately reflects the empirical probabilities across the 352 reduced elections than the partial-ballot IAC model. For other voting paradoxes such as monotonicity paradoxes or the spoiler effect, the IAC model typically gives probabilities much higher than what we observe in real elections. If we focus only on the 151 three-candidate actual elections, Dir(0.6)(0.6) gives the best approximation of the empirical unconditioned results (with the exception of “PL and RU”), which makes sense because in these elections, the winner is stronger overall than in the Dir(1)(1) (IAC) model, consistent with the election data. Overall, the 1D-spatial models, both normal and bimodal, underestimate the probabilities of one or more types of paradox occurring.
@@ 6.2. Comparison to Previous Reinforcement Paradox Results
As mentioned in the Introduction, we are aware of only two previous papers [12, 47] which estimate the probability with which an election demonstrates a reinforcement paradox. Our work differs from theirs in several significant ways. Most importantly, in most cases the analysis in these earlier papers was based on simulating two separate elections, checking whether the elections had the same winner, and then indicating a reinforcement paradox had occurred if the winner obtained from the combined election was different from the winner in the two generated elections. That is, the authors did not start with one election and look for a partition of the ballots as we do here. As Plassmann and Tideman state in [47], “We determine the frequencies of the reinforcement paradox by sampling ballots for two separate elections, which we then combine to determine whether the paradox occurred for any voting rule. In contrast, we determine the frequencies of all other paradoxes from a single sample of ballots.” Plassmann and Tideman do not try to decompose one set of ballots into a partition as we do, stating “There is no obvious way in which to divide the ballots from a single election into two distinct sets that would yield the appropriate distribution of rankings in the two new sets and would permit us to examine reinforcement.” Courtin et al. [12] generally take a similar approach, simulating two ballot sets and then combining them. We have contributed to the investigation of this paradox by identifying a method in the 3-candidate case, through Propositions 1 and 2, to determine if such a division is possible, without having to separately test each possible partition of a ballot set into two sub-elections. We feel our approach is equally valid since we are interested in the “paradoxical harm” felt by a candidate (and perhaps their supporters) in an election exhibiting the paradox: if there exists a partition of the electorate such that one of the losing candidates can point to each sub-electorate and claim victory, that candidate can reasonably claim some sort of harm when they ultimately lose the overall election.
@@
Finally, both previous papers report extremely small probabilities for observing a reinforcement paradox under IRV. Plassmann and Tideman give a probability of approximately 0.02% when using an electorate of at least 1,000 voters. Courtin et al. report similarly small probabilities. Using the term “Young’s reinforcement axiom” (YRA) synonymously with Young’s notion of consistency (Definition 1), Courtin et al. state, “Our results show that … cases of violations of YRA are rather rare… [[A]]lthough all these [[iterative scoring]] rules violate YRA, the frequencies of violation are very rare. Then, since the violation of this axiom is not so frequent, one should not be unduly worried about its theoretical possibility.” By contrast, we report extremely large reinforcement paradox probabilities, which are generally much larger than the probabilities of observing other classical paradoxes. The reason for this discrepancy is our difference in approach, as outlined above. We would obtain similarly small probabilities if, for example, we were to look at each real election and search for a reinforcement paradox by choosing a partition of the ballots at random. Our probabilities are higher because we consider all possible partitions simultaneously.
@@ 7. Conclusion
Both theoretical and empirical analysis suggests that 3-candidate IRV elections are highly susceptible to the reinforcement paradox. As noted above, this is partly because we consider all possible partitions. If we were require the two sub-elections to be similarly-sized, for instance, as they might be in two precincts, there will most likely be fewer instances of the paradox; the same is true if we restrict the sub-elections so that one sub-election is a small perturbation of the original election. Future work could examine how the probability of a reinforcement paradox changes under restrictions of this kind, or to instances involving a larger number of sub-elections.
Future work could also focus on geography-based decompositions of the electorate. Informally, study of the reinforcement paradox is often motivated in geographic terms using a story like this: Suppose a mayoral candidate wins when using only the ballots from the eastern half of a city and wins when using only the ballots from the western half. Then the candidate should win the overall election; a loss would be additively paradoxical. Some of our real-world data comes equipped with precinct information, allowing for geographical analysis. In practice, the number of precincts is often too large to check all possible decompositions of the electorate two precinct subsets, but perhaps heuristic algorithms could be developed to investigate if there are decompositions of the electorate into two sets of precincts, and if these sets of precincts are geographically clustered. The only such example of which we are aware is 2021 Ward 2 City Council IRV election in Minneapolis, in which candidate Yusra Arab wins when using only the ballots from precincts 1 and 5, and wins when using the ballots from all other precincts, but loses the overall election [35]. The number of precincts in Minneapolis’ second ward is small and this example was found using brute force.
Finally, we focus on IRV in this article because of its use in real-world political elections, but it would be interesting to repeat our analysis for other voting methods which are not positional scoring rules. For example, many researchers in the social choice space feel that Condorcet consistent methods are superior to IRV (see Chapter 2 of [13], for example), and thus we would welcome an examination of the susceptibility of Condorcet consistent methods to the reinforcement paradox.